Please help solve this problem. Thank You
Solve each system by the addition method. Determine whether the equations are independent, dependent, or inconsistent.
5x - y = 1
10x - 2y = 2

Answer by drglass(89)

There are two general ways to solve this problem, in this case the easiest is elimination. Given the two equations
5x - y = 1
10x - 2y = 2
we can multiply the top equation by 2 to get
10x - 2y = 2
This equation is identical to the second equation so the are dependent, you cannot solve this problem.